A 11 metre ladder is placed against a wall. The bottom of the ladder is 4m from the wall. If the ladder bottom slips to 4.5 how far down the wall does the top of the ladder slip. Ive tried doing it but each time i hand in the work it is wrong help please.
Answer by Boreal(15235)

Both are right triangles, with hypotenuse 11, the ladder.
The first is 4^2+x^2=11^2; 16 +x^2=121
x^2=105
x= +10.247 m
Now it is (4.5)^2+x^2=121;; 20.25 +x^2=121;
x^2=100.75
x=10.04 m
It would slip about 0.2 m.